# EventCast: Hybrid Demand Forecasting in E-Commerce with LLM-Based Event Knowledge

### v1

@@ Appendix C Appendix: Comparison with Native Time Series Models
EventCast consistently outperforms all baselines. On overall performance (Table 7), EventCast achieves an average MAE of 0.098 and MSE of 0.191, improving over the best native baseline (TimesNet, MAE = 0.122, MSE = 0.325) by 19.7% in MAE and 41.2% in MSE. During event-driven periods (Table 8), the gap widens dramatically: EventCast (MAE = 0.069, MSE = 0.089) outperforms TimesNet (MAE = 0.207, MSE = 0.752) by 66.7% in MAE and 88.2% in MSE.
Performance gap grows with forecast horizon and event volatility. At T+1, native models remain competitive. In Country 01 (event-driven), TimeMixer and OneFitsAll (MAE 0.035) slightly outperform EventCast (MAE 0.041), since short-horizon forecasts rely more on recent trends than future context. However, by T+4 the gap becomes substantial: in Country 01, EventCast achieves MAE of 0.033 vs. TimesNet’s 0.136 (75.7% reduction); in Country 02, MAE 0.087 vs. TimesNet’s 0.420 (79.3% reduction). This confirms that event knowledge becomes increasingly critical at longer horizons where historical patterns alone provide diminishing predictive power.
Countries with higher event volatility show the largest gains. Country 02 and Country 03 exhibit the most dramatic improvements during event-driven periods. In Country 03, EventCast achieves an average event-driven MAE of 0.128 compared to the best native baseline’s 0.325 (TimesNet) (60.6% reduction) reflecting the high event density and demand variability in this market. In contrast, Country 01 shows competitive performance even from native models at short horizons, consistent with its relatively more stable demand patterns.
Model architecture alone is insufficient for event-driven forecasting. The native baselines span diverse architectures: attention-based (PatchTST, iTransformer, PAttn), decomposition-based (Autoformer, DLinear), temporal convolution (TimesNet, TimeMixer), and LLM-based (Time-LLM, OneFitsAll). Despite architectural diversity, all native models suffer large errors during event-driven periods—the worst-performing models (DLinear, Autoformer) exhibit average event-driven MAE of 0.441 and 0.509, which are $6.4\times$ and $7.4\times$ that of EventCast. This confirms that the bottleneck is not model expressiveness but access to future event information.
TimesNet emerges as the strongest native baseline. TimesNet achieves the best native performance in both overall (MAE 0.122) and event-driven (MAE 0.207) settings, likely due to its multi-period temporal decomposition which partially captures recurring event patterns from historical data. Nevertheless, even TimesNet falls far short of EventCast, particularly at longer horizons and during volatile periods, confirming that learned periodicity cannot substitute for explicit event reasoning.

### v2

@@ Appendix C Appendix: Comparison with Native Time Series Models
EventCast consistently outperforms all baselines. On overall performance (Table 7), EventCast achieves an average MAE of 0.884 and MSE of 1.723, improving over the best native baseline (TimesNet, MAE = 1.098, MSE = 2.923) by 19.5% in MAE and 41.1% in MSE. During event-driven periods (Table 8), the gap widens dramatically: EventCast (MAE = 0.619, MSE = 0.801) outperforms TimesNet (MAE = 1.859, MSE = 6.767) by 66.7% in MAE and 88.2% in MSE.
Performance gap grows with forecast horizon and event volatility. At T+1, native models remain competitive. In Country 01 (event-driven), TimeMixer (MAE 0.315) and OneFitsAll (MAE 0.351) slightly outperform EventCast (MAE 0.368), since short-horizon forecasts rely more on recent trends than future context. However, by T+4 the gap becomes substantial: in Country 01, EventCast achieves MAE of 0.293 vs. TimesNet’s 1.228 (76.1% reduction); in Country 02, MAE 0.782 vs. TimesNet’s 3.779 (79.3% reduction). This confirms that event knowledge becomes increasingly critical at longer horizons where historical patterns alone provide diminishing predictive power.
Countries with higher event volatility show the largest gains. Country 02 and Country 03 exhibit the most dramatic improvements during event-driven periods. In Country 03, EventCast achieves an average event-driven MAE of 1.147 compared to the best native baseline’s 2.351 (OneFitsAll) (51.2% reduction) reflecting the high event density and demand variability in this market. In contrast, Country 01 shows competitive performance even from native models at short horizons, consistent with its relatively more stable demand patterns.
Model architecture alone is insufficient for event-driven forecasting. The native baselines span diverse architectures: attention-based (PatchTST, iTransformer, PAttn), decomposition-based (Autoformer, DLinear), temporal convolution (TimesNet, TimeMixer), and LLM-based (Time-LLM, OneFitsAll). Despite architectural diversity, all native models suffer large errors during event-driven periods—the worst-performing models (DLinear, Autoformer) exhibit average event-driven MAE of 3.966 and 4.578, which are $6.4\times$ and $7.4\times$ that of EventCast. This confirms that the bottleneck is not model expressiveness but access to future event information.
TimesNet emerges as the strongest native baseline. TimesNet achieves the best native performance in both overall (MAE 1.098) and event-driven (MAE 1.859) settings, likely due to its multi-period temporal decomposition which partially captures recurring event patterns from historical data. Nevertheless, even TimesNet falls far short of EventCast, particularly at longer horizons and during volatile periods, confirming that learned periodicity cannot substitute for explicit event reasoning.